OpenStudy (anonymous):

write an equation of the line containing the given point and perpendicular to the given line (7,-5); 3x+4y=7

OpenStudy (anonymous):

4y=-3x+7 y=-3/4x+7/4 slope =-3/4 perpendicular = -3/4 * x=-1 x=4/3 4/3(x,y)(7,-5) change in y --------- change in x y+5/x-7=4/3 3(y+5)=4(x-7) 3y+15=4x-28 3y=4x-13 y=4/3x-13/3